# ERX: A Fast Real-Time Anomaly Detection Algorithm for Hyperspectral Line-Scanning

### v1

@@ Abstract
Detecting unexpected objects (anomalies) in real-time has great potential for monitoring, managing, and protecting the environment. Hyperspectral line-scan cameras are a low-cost solution that enhance confidence in anomaly detection over RGB and multispectral imagery. However, real-time algorithms for these cameras must be fast when using small computers (e.g., those onboard a drone or small satellite), scalable to high dimensions, adaptable to changing scenery, and robust against geometric and radiometric distortions. This paper introduces the Exponentially moving RX algorithm (ERX) and compares it to existing RX-based anomaly detection methods for real-time line-scanning. ERX was tested using a Jetson Xavier NX compute module, achieving the best combination of speed and detection across three novel datasets compared to the other algorithms. This research paves the way for future studies in grouping and locating anomalous objects, adaptive and automatic threshold selection, and real-time field tests. The Python code for the algorithms and experiments is available at https://github.com/WiseGamgee/HyperAD.
@@
Anomaly Detection, Hyperspectral, Line-Scanning, Real-Time, Unsupervised Learning.
@@ I Introduction
Detecting anomalies in real-time offers an advantage for applications that require fast and autonomous decision making. This has great potential in hyperspectral imaging for remote sensing; a rapidly growing area of research used to monitor, manage, and protect the environment [5, 6, 7, 8, 9, 10]. Hyperspectral imaging, also known as imaging spectroscopy, provides greater confidence than RGB or multispectral imaging when detecting anomalies. This is because the pixels in a hyperspectral image (HSI) have more detailed spectral profiles that uniquely represent the material present [11], as shown in Figure 1. The greater depth of spectral information can make it easier to separate anomalies from the background [1].
This study focuses specifically on anomaly detection using hyperspectral line-scan cameras, as they are an increasingly popular choice of camera for remote sensing applications. Real-time anomaly detection applications that use these cameras include search and rescue [12], defence surveillance and reconnaissance [13], and hazard detection in agricultural harvesting [14].
Hyperspectral line-scan cameras capture one line of pixels at a time and require the motion of an attached platform such as a drone, aircraft, or satellite to collect an image (Figure 2). They have relatively high signal-to-noise ratios [15], high spatial-spectral resolutions [16], and are easier and cheaper to build with commercially available components.
However, line-scan cameras are sensitive to unexpected platform movement and vibrations, causing distortions that require geometric correction when creating an image [16]. Each line also needs radiometric correction to convert the raw line-scan data into physically meaningful units, such as reflectance. Both processes are challenging and time-consuming, and are typically performed in post-processing. Therefore, it is proposed that real-time anomaly detection algorithms should have the following characteristics:
  • Speed - fast processing for real-time results is essential to match the line-scan camera’s image capture, so algorithm compute time should be minimised for field implementation (e.g., onboard drones or satellites).
  • Scalability - leverages the numerous bands in HSIs.
  • Adaptability - manages concept drift and non-stationary behaviour, referring to changing distribution of imagery over time, crucial for monitoring large areas with varying scenery and lighting.
  • Robustness - maintains strong performance when operating with geospatial distortions present and less radiometric correction (e.g., using radiance instead of reflectance).
@@
The benchmark method for the detection of hyperspectral anomalies is the Reed-Xiaoli (RX) algorithm [17]. Given a hyperspectral image, the mean pixel vector (1) and the covariance matrix (2) are calculated using every pixel:
@@
where $\bm{x}_{i}\in\mathbb{R}^{b}$ represents each pixel, $b$ is the number of hyperspectral bands, and $n$ is the total number of pixels in the image. The mean and covariance model the image background using a normal distribution. The Mahalanobis distance (3) is then calculated for each pixel:
@@
The Mahalanobis distance gives an estimate of how far each pixel lies from the background, where a larger distance means a greater likelihood of being an anomaly. Finally, pixels are classified as anomalous if their distance is above a user-defined threshold ($T$):
$\hat{y}_{i}$ =$=\begin{cases}1,&\text{if }\delta_{i}\geq T\\ 0,&\text{otherwise}\end{cases}$ (4)
Although the RX algorithm is effective for identifying anomalies in HSIs, it is not designed for real-time implementation. Real-time implementation requires a sequential, time-series approach that only accesses past and present data. Many more methods have been established since the RX algorithm for detecting anomalies in HSIs, ranging from simpler traditional statistics and distance-based metrics to more complex unsupervised deep learning models [18]. However, only a small fraction of these anomaly detection algorithms are suitable for processing line-scan imagery in real-time. As such, the aims of this research are:
  1. 1.
    To provide a review of real-time RX-based hyperspectral line-scan anomaly detection algorithms.
  2. 2.
    To present the Exponentially moving RX anomaly detector (ERX); an algorithm that is fast, scalable, adaptable, and robust.
  3. 3.
    To evaluate ERX on a small edge computer and benchmark it’s performance against similar algorithms using various high-dimensional datasets.
II Related Work
This section summarises RX-based algorithms that are used for real-time line-scan anomaly detection. The algorithms are separated into subsections that share common methods. A limitation of the RX algorithm is that it requires capturing every single pixel before calculating the mean and covariance of a HSI. This prevents detecting hyperspectral anomalies in real-time and easily results in insufficient memory due to the large amount of data in HSIs. Furthermore, calculating the inverse of the covariance matrix is a slow operation in the RX algorithm because of the large number of hyperspectral bands. A common solution to these issues is to perform recursive mean and covariance updates, of which one popular method is the Woodbury Matrix Identity.
II-A The Woodbury Matrix Identity
The Woodbury matrix identity (5) is an expression that has been used to recursively update the inverse covariance matrix as new lines are captured from the camera, reducing both memory requirements and compute time.
@@ II-A The Woodbury Matrix Identity
Identifying $\bm{A}=\frac{n-1}{n}\bm{K}_{t-1}$, $\bm{u}=\bm{v}=\frac{1}{\sqrt{n}}(\bm{x}_{t}-\bm{\mu}_{t})$, and $c=1$, these values are substituted into the RHS of (5) and the inverse covariance matrix is updated.
@@
Furthermore, RX-BIL randomly removed pixels to increase detection speed. More pixels can be removed in homogeneous images as fewer are needed to accurately estimate the background mean and covariance. Fewer pixels should be removed in diverse images, so that more variable spectra are captured by the background statistics.
Liu et al. [21] combined the Woobury matrix identity with a finite Markov model for a line-wise anomaly detection algorithm, FMLRT-RAD. The addition of the finite Markov model improved the detection stability and adaptability by removing “redundant” information from previous lines. Compared to RT-CR-RXD, Liu et al. found that FMLRT-RAD had slightly higher detection performance and slightly slower, but comparable, processing speed.
II-B Cholesky Decomposition
While the Woodbury matrix identity offers a fast way to update the inverse covariance, Cholesky decomposition paired with forward substitution avoids calculating the inverse covariance altogether. Zhang et al. [22] proposed CDLSS, a line-wise anomaly detection algorithm that used Cholesky decomposition. Given a new line of pixels $\bm{X}_{t}$, the corresponding Mahalanobis distance of each pixel defined as:
@@ II-B Cholesky Decomposition
By substituting in $(\bm{X}_{t}\bm{L}_{t-1}^{-1})=\bm{z}_{i}$, the Mahalanobis distance is represented as:
$\delta_{i}$ $=\sqrt{\bm{z}_{i}\cdot\bm{z}_{i}}$ (11)
As $\bm{x}_{i}$ is known and $\bm{L}_{t-1}$ is calculated from Cholesky decomposition, forward substitution is used to calculate $\bm{z}_{i}$ directly, given that $\bm{z}_{i}\bm{L}_{t-1}=\bm{x}_{i}$.
Zhao and Xi-Feng [23] presented LRT-KRXD-CD, using Cholesky decomposition to reduce the processing time of the local real-time kernel RX detector (LRT-KRXD) [24]. Zhao et al. [25] presented RT-KCRD, building on KCRD [26] with Cholesky decomposition method. RT-KCRD used a moving window (discussed in the following section), recursive updates for the kernel covariance matrix, and an optimised regularisation matrix using previous anomaly scores. RT-KCRD was on average 86 times faster than KRXD and KCRD, and was comparable to LRT-KRXD-CD [25].
II-C Local/Window Algorithms
Window-based algorithms extract anomalies locally rather than at a global level, improving detection performance [27]. As shown in Figure 3, a standard window is a small region surrounding the current pixel being processed. Double-windows discard a small inner region of pixels, minimising the contribution of the same object to the local background statistics.
@@ II-C Local/Window Algorithms
Rossi et al. [28] improved two local anomaly detection algorithms using fast updates of the inverse covariance. The first, RX-LBL-IU, builds off the work of Acito et al. [29], which used “sliding” double-windows around each pixel to identify anomalies. The sliding window approach used the Woodbury matrix identity to recursively update the window inverse covariance, instead of recalculating it for each pixel’s window. The second method is Dark HORSE RX [13], which used an exponentially rolling mean and covariance for pixel-wise updates. Rossi et al. [28] deemed RX-LBL-IU the most effective of the two algorithms.
Zhao et al. [30] presented a causal pixel-wise approach, only accessing the window pixels captured before the current pixel being analysed. A two-step Woodbury matrix identity was used to update the inverse correlation matrix by adding the latest pixel and removing the oldest one.
Although window methods provide localised anomaly detection, a major challenge is calculating the window statistics efficiently. Small window sizes commonly led to matrix singularities, and large window sizes led to slow compute times. Preliminary experiments showed that they were too slow for real-time detection on the Jetson using Python.
II-D Kernel Methods
Kernel methods addressed the assumption of a normal distribution in the RX algorithm and subsequently improved detection accuracy. A kernel function allows the RX algorithm to map the non-linear correlations that hyperspectral bands share between one another in practice [31]. PLP-KRXD was proposed by Zhao et al. [32], using parallel sliding windows and the Woodbury matrix identity to speed up the KRX algorithm [31]. The sliding windows were implemented as a faster alternative to the double-window model. The Woodbury matrix identity reduced the time required to calculate the inverse covariance of multiple sliding windows. Zhao et al. demonstrated that PLP-KRXD achieved similar detection performance to KRXD, but was 40 to 58 times faster depending on the dataset. Kernel methods, despite their higher anomaly detection accuracy, are slower due to additional matrix operations.
II-E Dimensionality Reduction
Dimensionality reduction methods improve processing speed and signal-to-noise ratio by transforming hyperspectral bands into fewer features. Horstrand et al. [14] developed LbL-AD, which uses Principal Component Analysis (PCA) to reduce the number of bands. A line-wise subspace calculation using the power method and deflation was used to efficiently approximate the eigenvalues ${e}_{t}$ and the eigenvectors ${E}_{t}$. Using orthogonal subspace projection, the dimensions of the hyperspectral line were reduced:

### v2

# ERX: A Fast Real-Time Anomaly Detection Algorithm for Hyperspectral Line Scanning

@@ Abstract
Detecting unexpected objects (anomalies) in real time has great potential for monitoring, managing, and protecting the environment. Hyperspectral line-scan cameras are a low-cost solution that enhance confidence in anomaly detection over RGB and multispectral imagery. However, existing line-scan algorithms are too slow when using small computers (e.g. those onboard a drone or small satellite), do not adapt to changing scenery, or lack robustness against geometric distortions. This paper introduces the Exponentially moving RX algorithm (ERX) to address these issues, and compares it with existing RX-based anomaly detection methods for hyperspectral line scanning. Three large and more complex datasets are also introduced to better assess the practical challenges when using line-scan cameras (two hyperspectral and one multispectral). ERX is evaluated using a Jetson Xavier NX compute module, achieving the best combination of speed and detection performance. This research paves the way for future studies in grouping and locating anomalous objects, adaptive and automatic threshold selection, and real-time field tests. The datasets and the Python code are available at https://github.com/WiseGamgee/HyperAD.
@@
Anomaly Detection, Hyperspectral, Line Scanning, Real Time, Unsupervised Learning.
@@ I Introduction
Detecting anomalies in real time offers an advantage for applications that require fast and autonomous decision making. This has great potential in hyperspectral imaging for remote sensing; a rapidly growing area of research used to monitor, manage, and protect the environment [5, 6, 7, 8, 9, 10]. Hyperspectral imaging, also known as imaging spectroscopy, provides greater confidence than RGB or multispectral imaging when detecting anomalies. This is because the pixels in a hyperspectral image (HSI) have more detailed spectral profiles that uniquely represent the material present [11], as shown in Figure 1. The greater depth of spectral information can make it easier to separate anomalies from the background [1].
This study focuses specifically on anomaly detection using hyperspectral line-scan cameras, as they are an increasingly popular choice of camera for remote sensing applications. Line-scan cameras capture one line of pixels at a time and require the motion of an attached platform such as a drone, aircraft, or satellite to collect an image (Figure 2). They have relatively high signal-to-noise ratios [12], high spatial and spectral resolutions [13], and are easier and cheaper to build with commercially available components [14, 15]. Real-time anomaly detection applications that use these cameras include search and rescue [16], defence surveillance and reconnaissance [17], and hazard detection in agricultural harvesting [18].
@@
An essential consideration for line-scan cameras is their sensitivity to unexpected platform movements and vibrations, which causes distortions that require geometric correction when creating an image from a collection of lines [13]. Radiometric correction is also typically performed to convert raw data into physically meaningful units such as reflectance. However, geometric and radiometric corrections are complex and time-consuming, and are often unsuitable for real-time applications. Ideally, effective algorithms should function without these corrections.
Hyperspectral anomaly detection (HAD) has attracted significant research attention, with methods evolving from simpler traditional statistics and distance-based metrics to more complex unsupervised deep learning models [19]. The HAD benchmark is the Reed-Xiaoli (RX) algorithm [20], which calculates the Mahalanobis distance between each pixel and the background of a given HSI. The background is modelled as a normal distribution, represented by the mean and covariance of all pixels in this HSI.
In recent years, many new approaches have emerged. Collaborative representation-based methods use a regularised linear model to estimate a pixel from its neighbours, identifying anomalies by high prediction errors [21, 22, 23, 24]. Sparse representation adds an additional constraint by limiting the model to estimate the pixel using a few bands, reducing noise and improving anomaly detection [25, 26, 27]. Low-rank representation methods decompose the HSI into a low-rank background component, detecting anomalies that deviate from this representation [28, 29, 30]. Tensor-based methods have also been used to isolate anomalies by modelling the background, but preserve the spatial and spectral structure of the image [31, 32, 33]. This makes them well suited to the high dimensionality of hyperspectral data. Deep learning is an area of rapid growth, with various auto-encoding techniques used to reconstruct HSIs [34, 35, 36, 37, 38, 39]. Autoencoders typically detect anomalies through large reconstruction errors. Many additional works combine a variety of these different methods [40, 41, 42]. Some others even use chessboard topology [43], an “effective anomaly space” for improved anomaly-background separation [44], component decomposition [45], shallow graph neural networks [46], and more [47, 48].
However, none of the aforementioned algorithms are designed for real-time processing, let alone for the additional complexities of line scanning. Real-time line scanning is a causal time series process, accessing only the current line of pixels captured by the camera at time $t$ and previous lines [1, 49]. Most HAD methods analyse an image of the entire scene at once, and are only demonstrated on powerful CPUs and GPUs. This requirement for high computational and memory resources makes them unsuitable for applications onboard platforms with limited computing power. In general, real-time HAD is a less explored field that has an established need for more research [19, 50, 51, 52].
Variants of the RX algorithm have emerged as a popular choice for line-scan anomaly detection due to their effectiveness and lightweight design. They use different techniques to process the hyperspectral data pixel-by-pixel or line-by-line, making them ideal for real-time applications. One technique is the Woodbury matrix identity, which is used to recursively update the inverse correlation matrix Mahalanobis distance calculations [49, 53, 54]. Calculating the inverse is an otherwise slow and memory-intensive operation if performed for each new line. Other works apply Cholesky decomposition and forward substitution to avoid calculating the inverse covariance altogether [55, 56].
Window-based methods were used to improve anomaly detection by considering only the local surroundings of each pixel [57, 58, 59, 60, 61, 62], and kernel methods were found to enhance detection by better mapping the relationships between hyperspectral bands [63]. Algorithms using dimensionality reduction have also been proposed, using real-time versions of principal component analysis (PCA) and subspace projection [18, 64]. These reduce the number of spectral bands to improve the processing speed. However, even with these techniques, line scan anomaly detection algorithms lack a combination of three key characteristics:
  • Scalability - most algorithms struggle to handle high band numbers and line widths onboard small, low-power computing devices; a crucial requirement for real-time anomaly detection onboard drones or satellites.
  • Adaptability - many algorithms do not efficiently manage changing scenery or lighting, which is essential when monitoring large and diverse locations. This is known as concept drift or non-stationary behaviour, where the underlying distribution of an image changes over time.
  • Robustness - The literature has overlooked demonstrating algorithm stability and performance on larger datasets, particularly with the geometric distortions of line scanning. Previous studies often assessed algorithms on small, geometrically corrected datasets such as the Airport-Beach-Urban (ABU) dataset [65] or San Diego Airport, with only a few hundred pixels in height and width. Algorithms need testing on complex datasets to assess their suitability for real-time remote sensing applications.
To overcome these challenges, the Exponentially moving RX algorithm (ERX) is proposed. ERX processes hyperspectral images line-by-line, simultaneously updating the underlying background model and detecting anomalies. Sparse random projection (SRP) is used to reduce the dimensionality of the raw hyperspectral bands, allowing ERX to scale to high dimensions. ERX also uses exponentially moving averages (EMA) to rapidly update the mean and covariance, and quickly adapt to changing scenery by prioritising recent information. Cholesky decomposition is also used to further improve performance. ERX demonstrates robustness by outperforming similar algorithms on larger, more complex datasets that include geometric distortion and changing scenery. In summary, the overall contributions of this work are:
  1. 1.
    A thorough review of RX-based anomaly detection algorithms for line scanning. These algorithms are grouped by similar methods to provide an intuitive summary of the related work.
  2. 2.
    The Exponentially moving RX anomaly detector (ERX). ERX employs sparse random projection, exponential moving averages, and Cholesky decomposition to make it fast, scalable, adaptable, and robust. ERX outperforms similar algorithms in both speed and detection when evaluated on a small edge computer.
  3. 3.
    Three novel datasets for anomaly detection. These datasets are substantially larger than those used in the literature and include examples of changing scenery and geometric distortions. These datasets better capture the practical challenges of real-time anomaly detection for hyperspectral line scanning.
  4. 4.
    HyperAD: https://github.com/WiseGamgee/HyperAD, an openly available Python repository that contains the implementation of ERX and other algorithms from the literature. This is the first open-source repository for real-time anomaly detection for hyperspectral line scanning, which promotes accessibility and future research.
The remainder of this paper is structured as follows. Section II reviews related work on real-time anomaly detection for hyperspectral line scanning. Section III details the proposed method, ERX. Section IV covers the datasets and experimental design. Section V presents and discusses the results, outlining the core findings, limitations, and future directions. Finally, Section VI concludes the paper.
II Related Work
This section reviews RX-based algorithms that are used for real-time anomaly detection in line scanning. These algorithms are online, meaning that they detect anomalies while updating their models. They are separated into subsections that share common approaches.
II-A The RX Algorithm
The RX algorithm [20] detects anomalous pixels with spectral signatures that noticeably differ from the background of a given hyperspectral image, which is modelled as a normal distribution. The background is defined by the mean pixel vector (1) and covariance matrix (2), calculated using all pixels in the image:
@@ II-A The RX Algorithm
where $\bm{x}_{i}\in\mathbb{R}^{b}$ represents each pixel, $b$ is the number of hyperspectral bands, and $n$ is the total number of pixels in the image. The Mahalanobis distance (3) is then calculated for each pixel:
@@
The Mahalanobis distance gives an estimate of how far each pixel lies from the background, where a larger distance means a greater likelihood of being an anomaly. Finally, pixels are classified as anomalies if their distance is above a user-defined threshold ($\tau$):
$\hat{y}_{i}$ =$=\begin{cases}1,&\text{if }\delta_{i}\geq\tau\\ 0,&\text{otherwise}\end{cases}$ (4)
A core limitation of the base RX algorithm is that it requires the capture and storage of every single pixel before calculating the mean and covariance. This prevents real-time detection and easily results in insufficient memory due to the large amount of data. Calculating the inverse of the covariance matrix is also a slow operation because of the large number of hyperspectral bands. A common solution to these issues is to perform recursive mean and covariance updates, of which one popular method is the Woodbury Matrix Identity.
II-B The Woodbury Matrix Identity
The Woodbury matrix identity (5) is an expression that has been used to recursively update the inverse covariance matrix as new lines are captured from the camera, reducing memory usage and processing time.
@@ II-B The Woodbury Matrix Identity
Given that $\bm{A}=\frac{n-1}{n}\bm{K}_{t-1}$, $\bm{u}=\bm{v}=\frac{1}{\sqrt{n}}(\bm{x}_{t}-\bm{\mu}_{t})$, and $c=1$, these values are substituted into the RHS of (5) and the inverse covariance matrix is updated.
@@
RX-BIL randomly removed pixels to increase detection speed. More pixels can be removed in homogeneous images as fewer are needed to accurately estimate the background mean and covariance. Fewer pixels should be removed in diverse images, so that more variable spectra are captured by the background statistics.
Although the Woodbury matrix identity reduces computation time, each line equally contributes to the background estimation regardless of its collection time. Therefore, it does not inherently adapt to changing scenery and risks detecting false anomalies from redundant data. Liu et al. [54] partially addressed this issue using a combined Woobury matrix identity and finite Markov model for the line-by-line anomaly detection algorithm FMLRT-RAD. FMLRT-RAD improved stability and adaptability by using only recent lines, demonstrating slightly higher detection performance than RT-CR-RXD. However, its processing speed remained comparable, indicating a need for further speed improvements.
II-C Cholesky Decomposition
Cholesky decomposition paired with forward substitution avoids calculating the inverse covariance altogether, offering a faster alternative to the Woodbury matrix identity. Zhang et al. [66] proposed CDLSS, a line-by-line anomaly detection algorithm that used Cholesky decomposition. Given a new line of pixels $\bm{X}_{t}$, the corresponding Mahalanobis distance of each pixel defined as:
@@ II-C Cholesky Decomposition
By substituting in $(\bm{X}_{t}\bm{L}_{t-1}^{-1})=\bm{m}_{i}$, the Mahalanobis distance is represented as:
$\delta_{i}$ $=\sqrt{\bm{m}_{i}\cdot\bm{m}_{i}}$ (11)
As $\bm{x}_{i}$ is known and $\bm{L}_{t-1}$ is calculated from Cholesky decomposition, forward substitution is used to calculate $\bm{m}_{i}$ directly, given that $\bm{m}_{i}\bm{L}_{t-1}=\bm{x}_{i}$. However, using the correlation matrix increases the instability of the algorithm because of the higher risk of singularities, which can cause it to fail because it cannot compute the inverse.
Zhao and Xi-Feng [55] presented LRT-KRXD-CD, using Cholesky decomposition to reduce the processing time of the local real-time kernel RX detector (LRT-KRXD) [67]. Zhao et al. [56] presented RT-KCRD, building on KCRD [21] with Cholesky decomposition. RT-KCRD used a moving window (discussed in the following section), recursive updates for the kernel covariance matrix, and an optimised regularisation matrix using previous anomaly scores. RT-KCRD was on average 86 times faster than KRXD and KCRD, and was comparable to LRT-KRXD-CD [56]. Although these algorithms demonstrated the speed-up potential of Cholesky decomposition, they remain relatively slow due to the kernel function (discussed in Section II-E). Overall, Cholesky decomposition has yet to be used in a robust and fast algorithm.
II-D Local/Window Algorithms
Window-based algorithms extract anomalies locally rather than at a global level, improving detection performance [57] and adaptability. As shown in Figure 3, a standard window is a small region surrounding the current pixel being processed. Double-windows discard a small inner region of pixels, minimising the contribution of the same object to the local background statistics.
@@ II-D Local/Window Algorithms
Rossi et al. [58] improved two local anomaly detection algorithms using fast updates of the inverse covariance. The first, RX-LBL-IU, builds upon the work of Acito et al. [59], which used “sliding” double-windows around each pixel to identify anomalies. The sliding window approach used the Woodbury matrix identity to recursively update the window inverse covariance, instead of recalculating it for each pixel’s window. The second method is Dark HORSE RX [17], which used an exponentially rolling mean and covariance for pixel-wise updates. Rossi et al. [58] deemed RX-LBL-IU the most effective of the two algorithms.
Zhao et al. [60] presented a causal pixel-wise approach, only accessing the window pixels captured before the current pixel being analysed. A two-step Woodbury matrix identity was used to update the inverse correlation matrix by adding the latest pixel and removing the oldest one. Liu et al. introduced two more local anomaly detection algorithms; one designed around multi-line multi-band processing [61] and another for pixel-by-pixel processing [62]. These algorithms update the correlation matrix recursively and focus on recent pixels.
However, calculating the window statistics efficiently is a major challenge. Small window sizes can lead to matrix singularities, and large window sizes lead to slow compute times. Preliminary experiments with Python were too slow for real-time detection, failing to meet the speed requirements. Many window approaches also assume geometrically corrected pixels, which can reduce the robustness for real-time line scanning.
II-E Kernel Methods
Kernel methods address the assumption of a normal distribution in the RX algorithm and subsequently improve detection accuracy. A kernel function allows the RX algorithm to map the non-linear correlations that hyperspectral bands share with each other in practice [68]. PLP-KRXD was proposed by Zhao et al. [63], using parallel sliding windows and the Woodbury matrix identity to speed up the KRX algorithm [68]. The sliding windows were implemented as a faster alternative to the double-window model. The Woodbury matrix identity reduced the time required to calculate the inverse covariance of multiple sliding windows. Zhao et al. demonstrated that PLP-KRXD achieved detection performance similar to KRXD, but was 40 to 58 times faster, depending on the dataset.
Kernel methods, though more accurate, are much slower due to the additional matrix operations required. This work focuses on non-kernel methods which are better suited to real-time processing with limited compute resources.
II-F Dimensionality Reduction
Dimensionality reduction methods improve processing speed and signal-to-noise ratio by transforming hyperspectral bands into fewer features. Horstrand et al. [18] developed LbL-AD, which uses Principal Component Analysis (PCA) to reduce the number of bands. Line-by-line subspace calculation using the power method and deflation was used to efficiently approximate the eigenvalues $\bm{e}_{t}$ and the eigenvectors $\bm{E}_{t}$. Using orthogonal subspace projection, the dimensions of the hyperspectral line were reduced: